I need help with this problem. Premium coffee sells for $6.00 per pound and regular coffee sells for $4.00 per pound. How many pounds of each type of coffee should be blended to obtain 100 pounds of a blend that sells for $4.64 per pound.
Heres what I was doing but I am getting confused.
6(100) + 4x = 4.64(100-x)
600 + 4x =4.64 - 4.64x
8.64x= -136
I think I have messed up somewhere.

Answer by Earlsdon(6294)

Ok, let x = the number of lbs of premium coffee at $6 per lb. Then (100 - x) = the number of lbs of regular coffee at $4 per lb. So you need the sum of these to make 100 lbs of blended coffee at $4.64 per lb.
Here's the equation:
x($6) + (100-x)($4) = 100($4.64) Simplify and solve for x, the number of lbs of premium coffee.
6x + 400 - 4x = 464
2x + 400 = 464 Subtract 400 from both sides.
2x = 64 Divide both sides by 2
x = 32 lbs of premium coffee.
100 - x = 100 - 32 = 68 lbs of regular coffee.
Check:
32(6) + 68(4) = 100(4.64)
192 + 272 = 464
464 = 464 Ok.